# Weyl law for 1-cycles

### v1

@@ 4.1. Parametric isoperimetric inequality in $\mathbb{D}^{n}$
Given $\varepsilon>0$, there is a piecewise linear map $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with the following properties. $\Psi$ is linear on each simplex of a certain triangulation of $\mathbb{R}^{n}$. Each top-dimensional simplex of this triangulation is labelled good or bad. For each good simplex $\Delta$, $\Psi(\Delta)$ lies in $S$. Each bad simplex lies in $N_{\varepsilon}T$. There exists a constant $C(n)$ depending only on $n$ such that the triangulation and the map obey the following bounds
@@
Let $1\leq j\leq n$. First we will show that the number of bad simplices $\Delta$ such that $\operatorname{spt}(\tau_{j})\cap\Delta\neq\emptyset$ is bounded by (${n\choose 2}C(n)$. As all bad simplices are contained in $N_{\varepsilon}T$ we see that for such $\Delta$, $\operatorname{spt}(\tau_{j})\cap\Delta$ is contained in $L_{j}\cap N_{\varepsilon}T$. By Lemma 4.2, $L_{j}\cap N_{\varepsilon}T$ is contained in the union of at most (${n\choose 2}$ balls $\{B_{i}^{L_{j}}\}$ of radius smaller than $1$. By property (1) in Lemma 4.4, the collection $\{B_{i}^{j}\}_{i}$ intersects at most (${n\choose 2}C(n)$ different bad simplices. Next we observe that if $\Delta$ is any simplex, by (2) and (3) in Lemma 4.4 we have $\operatorname{diam}(\Psi(\Delta))\leq 3C(n)$. Therefore as $\Psi(\tau_{j}\llcorner\Delta)$ is a linear segment,
@@
Consider a grid of width $r$ in the unit disk $\mathbb{D}^{n}$ with $1$-skeleton $S$ and dual $(n-2)$-skeleton $T$. Set $R=\frac{1}{r}$. Identify $\mathbb{D}^{n}$ with $B_{R}(0)$ under the map $x\mapsto Rx$ which maps the skeleta $S$ and $T$ of $\mathbb{D}^{n}$ to the corresponding $S$ and $T$ of $B_{R}(0)$ defined in Lemma 4.2. Use that lemma to pick a suitable point $P$ outside $\mathbb{D}^{n}$ and construct a continuous family of $1$-chains $G_{1}:X\to\mathcal{I}_{1}(\mathbb{D}^{n})$ in the following way. Given two points $z,w\in\mathbb{D}^{n}$ let $\Xi(z,w)$ be the segment connecting them. Given a 00-chain $x$ in $\mathbb{D}^{n}$ and a Lipschitz map $T:\mathbb{D}^{n}\to\mathbb{D}^{n}$ define
@@
Notice that $\partial G_{1}(x)=H(F(x))-F(x)=H(\overline{F}(x))-\overline{F}(x)$. Now we want to deform $G_{1}$ to decrease its mass, using a bend and cancel argument as in [11]. We use the previously constructed map $\Psi$. By Lemma4.5,
@@ 4.2. Two important propositions to go from disks to manifolds
To prove the Weyl law, we will need to deal with families of relative 00-cycles supported in a compact manifold with boundary $(M,\partial M)$ rather than with absolute families supported in $\mathbb{D}^{n}$ as in Theorem 4.1. To extend our result to such families, the strategy will be to pick a triangulation $\{Q_{j}\}_{j}$ of the underlying compact manifold $M$ where our family $F$ is supported and find a short filling of $F(x)\llcorner Q_{j}$ for each $j$. But $F(x)\llcorner Q_{j}$ is not a continuous family of absolute cycles in $Q_{j}$ but only a continuous family of relative cycles in $(Q_{j},\partial Q_{j})$ (even if $\partial M=\emptyset$). Assuming that the original family $F$ is contractible, it is possible to represent $F\llcorner Q_{j}$ by a family $F_{j}$ of absolute cycles. However, $F_{j}$ may have very large mass in $\partial Q_{j}$, even in arbitrary tiny balls in $\partial Q_{j}$ as the ones used in the proof of Theorem 4.1; preventing from its direct application to find a short filling of $F_{j}$. The next proposition provides a way to slightly perturb a family $F:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{n})$ into a new family $F^{\prime}(x)$ for which there exists a small ball $B\subseteq\mathbb{S}^{n-1}$ such that $\operatorname{\mathbf{M}}(F^{\prime}(x)\llcorner B)$ is controlled for all $x\in X$. We first state and prove the proposition for $n=3$ and then we discuss the case $n\geq 4$.
@@
Let $F:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{3})$ be a $\delta$-localized family. Let $L>0$ be a very small real number and let $B$ be the open ball in $\mathbb{S}^{2}$ centered at $e$ of radius $L$. There exists a continuous family $F^{\prime}:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{3})$ with the following properties:
@@
    If $\tilde{B}=B_{L+(p+2)\delta}(e)$ then $F^{\prime}(x)\llcorner(\mathbb{D}^{3}\setminus\tilde{B})=F(x)\llcorner(\mathbb{D}^{3}\setminus\tilde{B})$ for every $x\in X^{p}$.
@@
Before proving the proposition, we introduce the following definitions.
@@
Take polar coordinates $(r,\theta)\in[0,\infty)\times[0,2\pi)$ in the $xy$-plane and consider the induced cylindrical coordinates $(r,\theta,z)$ and spherical coordinates $(r,\theta,\varphi)$ in $\mathbb{R}^{3}$ ($z\in\mathbb{R}$, $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$). Given $\theta\in[0,2\pi)$, let $s_{\theta}$ denote the geodesic segment contained in the meridian passing by $e=(0,0,-1)$ and the point $(\cos\theta,\sin\theta,0)$ given by the set of points there with
$-\frac{\pi}{2}+\frac{L}{2}\leq\varphi\leq 0$
@@
$-\cos(\frac{L}{2})\leq z\leq 0.$
Given $10<\eta<$ 1 denote
@@
We define $F^{\prime}$ inductively skeleton by skeleton. At the 00-skeleton, we proceed as in [12]. Namely, given $x\in X_{0}$ we define
@@
where $e(x)=e$ (the 00-chain associated to the point $e\in\mathbb{S}^{2}$) if $\operatorname{\mathbf{M}}(F(x)\llcorner B)$ is odd and $e(x)=0$ otherwise. Notice that (2) and (3) hold and also that
@@
By Lemma 3.7, we can assume that $F$ is monotonously $\delta$-localized. Given $C$ a top dimensional cell of $X$, let $\{B_{i}^{C}\}_{i\in I_{C}}$ be the $\delta$-admissible family where $F|_{C}$ is localized. Let
@@
Given $C\in\operatorname{Faces}_{p}(X)$, let $R_{C}\in[L+\delta,L+2\delta]$ be a real number such that $\partial B_{R_{C}}(e)\cap B_{i}^{C}=\emptyset$ for every $i\in I_{C}$ and $\partial B_{R_{C}}(e)\cap\mathcal{S}=\emptyset$ (such $R_{C}$ exists because $\sum_{i\in I_{C}}\operatorname{diam}(B_{i}^{C})<\delta$). Denote
$J_{C}=\{i\in I_{C}:B_{i}^{C}\cap B_{R_{C}}(e)=\emptyset\}.$
We claim that there exists $\theta_{C}\in[0,2\pi)$ such that $s_{\theta_{C}}\cap B_{i}^{C}=\emptyset$ for every $i\in J_{C}$ and $s_{\theta}\cap\mathcal{S}=\emptyset$. Given $i\in J_{C}$, denote
@@
be the interval in the circle $\partial B_{R_{C}}(e)\cap\mathbb{S}^{2}$ consisting of those points with $\theta\in\tilde{I}_{i}$ in cylindrical coordinates. Notice that
@@
and as the length of the circle $\partial B_{R_{C}}(e)\cap\mathbb{S}^{2}$ is $2\pi\sin(R_{C})$ we can see
@@
provided $\delta$ is sufficiently small. Therefore, there must exist a point in $\partial B_{R_{C}}(e)\cap\mathbb{S}^{2}\setminus(\mathcal{S}^{\prime}\cup\bigcup_{i\in J_{C}}I_{i})$ which yields the existence of the desired $s_{\theta_{C}}$ (here $\mathcal{S^{\prime}}$ is the set of points in $\partial B_{R_{C}}(e)\cap\mathbb{S}^{2}$ whose $\theta$-coordinate coincides with that of some point in $\mathcal{S}$). By compactness, there exists $\eta_{C}>0$ such that the $\eta_{C}$-tubular neighborhood of $s_{\theta_{C}}$ in $\mathbb{D}^{3}$ does not intersect $\mathcal{S}\cup\bigcup_{i\in J_{C}}B_{i}^{C}$. We set $T_{C}=T_{\theta_{C},\eta_{C}}$. Then by construction, we have an $L+3\delta$-admissible family
$\{B_{i}^{C}\}_{i\in I^{*}_{C}}=\{B_{R_{C}}(e)\}\cup\{B_{i}^{C}:i\in J_{C}\}$
@@
$F^{\prime}(x)-F^{\prime}(y)=F(x)-F(y)-(F(x)-F(y))\llcorner B+e(x)-e(y)$
being $F(x)-F(y)$ supported in $\bigcup_{i\in I_{C}}B_{i}^{C}\subseteq\bigcup_{i\in I^{*}_{C}}B_{i}^{C}$ and
$\operatorname{spt}((F(x)-F(y))\llcorner B+e(x)-e(y))\subseteq B\subseteq B_{R_{C}}(e)\subseteq\bigcup_{i\in I^{*}_{C}}B_{i}^{C}.$
@@
  1. (D1)
    $F^{\prime}|_{E}$ is localized in $\{U_{i}^{E}\}_{i\in I_{E}^{*}}$.
  2. (D2)
    $F^{\prime}(x)\llcorner(\mathbb{D}^{3}\setminus B_{r_{E}}(e))=F(x)\llcorner(\mathbb{D}^{3}\setminus B_{r_{E}}(e))$.
  3. (D3)
@@
  4. (D4)
    $U_{i}^{E}\cap T_{C}=\emptyset$ for every $i\in J_{E}$ and every $C\in\operatorname{Faces}_{p}(X)$ containing $E$.
  5. (D5)
    For every $x\in E$
@@
Observe that $L+\delta\leq r_{E}\leq L+2\delta$ and that if $C_{0}\in\operatorname{Faces}_{p}(X)$ is such that $r_{E}=R_{C_{0}}$, then $\partial B_{r_{E}}(e)$ does not intersect $U_{i}^{C_{0}}$ for every $i\in I_{C_{0}}$ and hence neither intersects $U_{i}^{E}$ for any $i\in I_{E}$ (by monotonously $\delta$-localization). In addition, $\partial B_{r_{E}}(e)\cap\mathcal{S}=\emptyset$. Define $J_{E}$ and $\{U_{i}^{E}\}_{i\in I_{E}^{*}}$ as in the inductive property, being
$\bigcup_{i\in I_{E}}U_{i}^{E}\subseteq\bigcup_{i\in I_{E}^{*}}U_{i}^{E}.$
@@
$F^{\prime}(t)=F(t)\llcorner(\mathbb{D}^{3}\setminus B_{r_{E}}(e))+F_{E}(t)$
where we define $F_{E}$ as follows. As $\operatorname{\mathbf{M}}(F^{\prime}(x)\llcorner B_{r_{E}}(e))$ and $\operatorname{\mathbf{M}}(F^{\prime}(y)\llcorner B_{r_{E}}(e))$ have the same parity, we can homotop both $F^{\prime}(x)\llcorner B_{r_{E}}(e)$ and $F^{\prime}(y)\llcorner B_{r_{E}}(e)$ to $e(x)$ defined as 00 if $\operatorname{\mathbf{M}}(F^{\prime}(x)\llcorner B_{r_{E}}(e))$ is even and $e$ otherwise. To be precise, for $t\in[0,\frac{1}{2}]$ we construct $F_{E}(t)$ by moving one point $z\in\operatorname{spt}(F^{\prime}(x)\llcorner B_{r_{E}}(e))$ towards $e$ at a time. The trajectories $\gamma_{z}$ of such points are curves in $\mathbb{D}^{3}$ which do not intersect
@@
and which are supported in $\partial\mathbb{D}^{3}$ for $z\in\partial\mathbb{D}^{3}$. For $t=[\frac{1}{2},1]$ we proceed in the same way but starting from $F_{E}(1)=F^{\prime}(y)\llcorner B_{r_{E}}(e)$ and going towards $F_{E}(\frac{1}{2})=e(x)$. By construction it holds
@@
This yields (D5).
Inductive step. Let $E$ be a $(k+1)$-face of $X$. Choose $r_{E}\in[L+(k+1)\delta,L+(k+2)\delta]$ so that $\partial B_{r_{E}}(e)\cap U_{i}^{E}=\emptyset$ and $\partial B_{r_{E}}(e)\cap\mathcal{S}=\emptyset$ for every $i\in I_{E}$. Define $J_{E}$ and $\{U_{i}^{E}\}_{i\in I_{E}^{*}}$ as in the inductive property. Notice that $F^{\prime}|_{\partial E}$ is localized in $\{U_{i}^{E}\}_{i\in I_{E}^{*}}$ because of the monotonously $\delta$-localization of $F$ and the fact that

### v2

@@ 4.1. Parametric isoperimetric inequality in $\mathbb{D}^{n}$
Given $\varepsilon>0$, there is a piecewise linear map $\Psi:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with the following properties. $\Psi$ is linear on each simplex of a certain triangulation of $\mathbb{R}^{n}$. Each top-dimensional simplex of this triangulation is labeled good or bad. For each good simplex $\Delta$, $\Psi(\Delta)$ lies in $S$. Each bad simplex lies in $N_{\varepsilon}T$. There exists a constant $C(n)$ depending only on $n$ such that the triangulation and the map obey the following bounds
@@
Let $1\leq j\leq n$. First, we will show that the number of bad simplices $\Delta$ such that $\operatorname{spt}(\tau_{j})\cap\Delta\neq\emptyset$ is bounded by ${n\choose 2}C(n)$. As all bad simplices are contained in $N_{\varepsilon}T$ we see that for such $\Delta$, $\operatorname{spt}(\tau_{j})\cap\Delta$ is contained in $L_{j}\cap N_{\varepsilon}T$. By Lemma 4.2, $L_{j}\cap N_{\varepsilon}T$ is contained in the union of at most $n\choose 2$} balls $\{B_{i}^{L_{j}}\}$ of radius smaller than $1$. By property (1) in Lemma 4.4, the collection $\{B_{i}^{j}\}_{i}$ intersects at most ${n\choose 2}C(n)$ different bad simplices. Next we observe that if $\Delta$ is any simplex, by (2) and (3) in Lemma 4.4 we have $\operatorname{diam}(\Psi(\Delta))\leq 3C(n)$. Therefore as $\Psi(\tau_{j}\llcorner\Delta)$ is a linear segment,
@@
Consider a grid of width $r$ in the unit disk $\mathbb{D}^{n}$ with $1$-skeleton $S$ and dual (n2)(n-2)-skeleton $T$. Set $R=\frac{1}{r}$. Identify $\mathbb{D}^{n}$ with $B_{R}(0)$ under the map $x\mapsto Rx$ which maps the skeleta $S$ and $T$ of $\mathbb{D}^{n}$ to the corresponding $S$ and $T$ of $B_{R}(0)$ defined in Lemma 4.2. Use that lemma to pick a suitable point $P$ outside $\mathbb{D}^{n}$ and construct a continuous family of $1$-chains $G_{1}:X\to\mathcal{I}_{1}(\mathbb{D}^{n})$ in the following way. Given two points $z,w\in\mathbb{D}^{n}$ let $\Xi(z,w)$ be the segment connecting them. Given a 0-chain $x$ in $\mathbb{D}^{n}$ and a Lipschitz map $T:\mathbb{D}^{n}\to\mathbb{D}^{n}$ define
@@
Notice that $\partial G_{1}(x)=H(F(x))-F(x)=H(\overline{F}(x))-\overline{F}(x)$. Now we want to deform $G_{1}$ to decrease its mass, using a bend and cancel argument as in [GuthMinMax] and [GL22]. We use the previously constructed map $\Psi$. By Lemma 4.5,
@@ 4.2. Two important propositions to go from disks to manifolds
To prove the Weyl law, we will need to deal with families of relative 0-cycles supported in a compact manifold with boundary $(M,\partial M)$ rather than with absolute families supported in $\mathbb{D}^{n}$ as in Theorem 4.1. To extend our result to such families, the strategy will be to pick a triangulation $\{Q_{j}\}_{j}$ of the underlying compact manifold $M$ where our family $F$ is supported and find a short filling of $F(x)\llcorner Q_{j}$ for each $j$. But $F(x)\llcorner Q_{j}$ is not a continuous family of absolute cycles in $Q_{j}$ but only a continuous family of relative cycles in $(Q_{j},\partial Q_{j})$ (even if $\partial M=\emptyset$). Assuming that the original family $F$ is contractible, it is possible to represent $F\llcorner Q_{j}$ by a continuous family $F_{j}$ of absolute cycles. However, $F_{j}$ may have very large mass in $\partial Q_{j}$, even in arbitrary tiny balls in $\partial Q_{j}$ as the ones used in the proof of Theorem 4.1; preventing from its direct application to find a short filling of $F_{j}$. The next proposition provides a way to slightly perturb a family $F:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{n})$ into a new family $F^{\prime}(x)$ for which there exists a small ball $B\subseteq\mathbb{S}^{n-1}$ such that $\operatorname{\mathbf{M}}(F^{\prime}(x)\llcorner B)$ is controlled for all $x\in X$. We first state and prove the proposition for $n=3$ and then we discuss the case $n\geq 4$. The analog proposition for $n=2$ was proved by Guth and Liokumovich in [GL22][Lemma 5.2] and it was used to obtain the Parametric Isoperimetric Inequality for 0-cycles on surfaces.
@@
There exists $\delta_{0}=\delta_{0}(p)>0$ such that the following is true. Let $F:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{3})$ be a $\delta$-localized family, $\delta<\delta_{0}$. Let $0<L<\delta$ be a very small real number and let $B$ be the open ball in $\mathbb{S}^{2}$ centered at $e$ of radius $L$. There exists a continuous family $F^{\prime}:X^{p}\to\mathcal{Z}_{0}(\mathbb{D}^{3})$ with the following properties:
@@
[…]
@@
Before proving the proposition, we introduce the following definition.
@@
Take polar coordinates $(r,\theta)\in[0,\infty)\times[0,2\pi)$ in the $xy$-plane and consider the induced cylindrical coordinates $(r,\theta,z)$ and spherical coordinates $(r,\theta,\varphi)$ in $\mathbb{R}^{3}$ ($z\in\mathbb{R}$, $\varphi\in[-\frac{\pi}{2},\frac{\pi}{2})$). Given $\theta\in[0,2\pi)$, let $s_{\theta}$ denote the geodesic segment in $\mathbb{S}^{2}$ contained in the meridian passing by e=(0,0,1)e=(0,0,-1) and the point $(\cos\theta,\sin\theta,0)$ given by the set of points there with
$\frac{\pi}{2}+\frac{L}{2}\leq\varphi\leq 0$
@@
$\cos(\frac{L}{2})\leq z\leq 0.$
Given $0<\eta<1$ denote
@@
We define $F^{\prime}$ inductively skeleton by skeleton. At the 0-skeleton, we proceed as in [GL22]. Namely, given $x\in X_{0}$ we define
@@
where $e(x)=e$ (the 0-chain associated to the point $e=(0,0,-1)\in\mathbb{S}^{2}$) if $\operatorname{\mathbf{M}}(F(x)\llcorner B)$ is odd and $e(x)=0$ otherwise. Notice that (1) to (3) hold and also that
@@
To motivate our construction to extend $F^{\prime}$ to the higher-dimensional skeleta, let us first explain what is done in [GL22] for the case $n=2$. We first introduce the following notation. Given $r\in(0,\frac{\pi}{2})$, let
$B_{r}$ $=\{q\in\mathbb{S}^{1}:\operatorname{dist}_{\mathbb{S}^{1}}(q,e)\leq r\}$
$C_{r}$ $=\{(y,z)\in\mathbb{D}^{2}:z=-\cos(r)\}$
$U_{r}$ $=\{(y,z)\in\mathbb{D}^{2}:z\leq-\cos(r)\}.$
Observe that $\partial U_{r}=B_{r}\cup C_{r}$ and $B=B_{L}$. Denote $U=U_{L}$. Suppose $F^{\prime}$ was defined in $\partial C$ for a certain cell $C$ of $X$ verifying that $F^{\prime}|_{\partial C}$ is localized in the $\delta$-admissible family $\{U_{i}^{C}\}_{i\in I_{C}}$. Our goal is to contract $F^{\prime}|_{\partial C}$ in order to extend $F^{\prime}$ to all of $C$. Consider a radius $r_{C}\in[L,L+\delta]$ such that $F^{\prime}(x)\llcorner C_{r_{C}}=0$ for every $x\in\partial C$ (such $r_{C}$ exists by the $\delta$-localization property). Denote $U_{C}=U_{r_{C}}$ and $B_{C}=B_{r_{C}}$. The previous allows us to write $F^{\prime}|_{\partial C}$ as the sum of two continuous families
$F^{\prime}(x)=F^{\prime}(x)\llcorner(\mathbb{D}^{2}\setminus U_{C})+F_{C}(x)$
where $F_{C}(x)=F(x)\llcorner U_{C}$. The challenging part is how to contract $F_{C}$ without introducing too many points in $B$. In particular, just applying a radial homotopy on $U_{C}$ towards $e$ will not work, because there might be arbitrarily many points in $B_{C}\setminus B$ which are pushed to $B$ by this homotopy. Guth and Liokumovich proposed the following solution. Let $R_{C}:U_{C}\times[0,1]\to U_{C}$ be the map which is the identity on $U$ and retracts the region $U_{C}\setminus U$ radially onto the line $C_{L}$, mapping points in $\mathbb{S}^{1}$ to $\mathbb{S}^{1}$. For $x\in\partial C$, consider
$F_{C}(x,t)=R_{C}(F_{C}(x),t)$
Then it is clear that $F_{C}(x,t)\llcorner B=F_{C}(x)\llcorner B$ for every $t\in[0,1)$, and at $t=1$ because of $\mathbb{Z}_{2}$ cancellation we introduce at most two new points to $B$ (the only two points in $\mathbb{S}^{1}$ at distance $L$ from $e$). If we then homotop $F_{C}(x,1)$ to the constant family $F_{C}(x,2)=e$ by contracting $U$ radially towards $e$, we get an extension with the desired properties.
We can perform the same constructions on $\mathbb{D}^{3}$. In this case, the definitions are
$B_{r}$ $=\{q\in\mathbb{S}^{2}:\operatorname{dist}_{\mathbb{S}^{2}}(q,e)\leq r\}$
$C_{r}$ $=\{(x,y,z)\in\mathbb{D}^{3}:z=-\cos(r)\}$
$U_{r}$ $=\{(x,y,z)\in\mathbb{D}^{3}:z\leq-\cos(r)\}.$
Nevertheless, in this case $F_{C}(x,1)$ can potentially have arbitrarily many points in $\partial B$, because it is a circle (and not a set of two points as before). Therefore for $t\in(1,2)$, $F_{C}(x,t)\llcorner B$ can have a very large number of points. To fix the previous, we will perform a different homotopy to make sure that $F_{C}(x,1)\llcorner\partial B$ is supported in a discrete set which is independent of $x$, instead of in a $1$-dimensional set as we obtained with the previous construction. This will be obtained by first homotoping $F_{C}=F_{C}^{0}$ to a map $F^{1}_{C}$ for which $F^{1}_{C}\llcorner(B_{C}\setminus B)$ is supported in a finite collection of meridians and then applying the homotopy $R_{C}$ to $F^{1}_{C}$. To illustrate this, suppose that we want $F^{1}_{C}\llcorner(B_{C}\setminus B)$ to be supported in exactly two meridians $s_{\theta_{1}}$ and $s_{\theta_{2}}$, for example with $\theta_{1}=0$ and $\theta_{2}=\pi$. If we had two other meridians $s_{\theta_{1}^{\prime}}$ and $s_{\theta_{2}^{\prime}}$ (say $\theta_{1}^{\prime}=\frac{\pi}{2}$ and $\theta_{2}^{\prime}=\frac{3\pi}{2})$ such that $\operatorname{spt}(F_{C}\llcorner(B_{C}\setminus B))$ does not intersect $s_{\theta_{1}^{\prime}}\cup s_{\theta_{2}^{\prime}}$, then as $(B_{C}\setminus B)\setminus(s_{\theta_{1}^{\prime}}\cup s_{\theta_{2}^{\prime}})$ retracts onto $s_{\theta_{1}}\cup s_{\theta_{2}}$, we could obtain the desired $F^{1}_{C}$ by applying a suitable extension (to $U_{E}$) of the previous retraction to $F_{C}$. We will extend $F^{\prime}$ inductively such that for each cell $C$ of $X$, $F_{C}$ avoids a certain collection of meridians which depend on $C$, like $s_{\theta^{\prime}_{1}}$ and $s_{\theta^{\prime}_{2}}$ in the previous example. The latter property will allow us to extend $F^{\prime}$ to the skeleton which is $1$-dimension higher.
We start by explaining the construction of the meridians and then we proceed to build $F^{\prime}$ by induction. By Lemma 3.7, we can assume that $F$ is monotonously $\delta$-localized. Given $C$ a top dimensional cell of $X$, let $\{B_{i}^{C}\}_{i\in I_{C}}$ be the $\delta$-admissible family where $F|_{C}$ is localized. Let
@@
Given $C\in\operatorname{Faces}_{p}(X)$, let $r_{C}\in[L+\delta,L+2\delta]$ be a real number such that $C_{r_{C}}\cap B_{i}^{C}=\emptyset$ for every $i\in I_{C}$ and $C_{r_{C}}\cap\mathcal{S}=\emptyset$ (such $r_{C}$ exists because $\sum_{i\in I_{C}}\operatorname{diam}(B_{i}^{C})<\delta$). Denote
$J_{C}=\{i\in I_{C}:B_{i}^{C}\cap U_{r_{C}}=\emptyset\}.$
We claim that there exists $\theta_{C}\in[0,2\pi)$ such that $s_{\theta_{C}}\cap B_{i}^{C}=\emptyset$ for every $i\in J_{C}$ and $s_{\theta_{C}}\cap\mathcal{S}=\emptyset$. Given $i\in J_{C}$, denote
@@
be the interval in the circle $\partial B_{r_{C}}\subseteq\mathbb{S}^{2}$ consisting of those points with $\theta\in\tilde{I}_{i}$ in cylindrical coordinates. Notice that
@@
and as the length of the circle $\partial B_{r_{C}}$ is $2\pi\sin(r_{C})$ we can see
@@
provided $\delta$ is sufficiently small. Therefore, there must exist a point in $\partial B_{r_{C}}\setminus(\mathcal{S}^{\prime}\cup\bigcup_{i\in J_{C}}I_{i})$ which yields the existence of the desired $s_{\theta_{C}}$ (here $\mathcal{S^{\prime}}$ is the set of points in $\partial B_{r_{C}}$ whose $\theta$-coordinate coincides with that of some point in $\mathcal{S}$). By compactness, there exists $\eta_{C}>0$ such that the $\eta_{C}$-tubular neighborhood of $s_{\theta_{C}}$ in $\mathbb{D}^{3}$ does not intersect $\mathcal{S}\cup\bigcup_{i\in J_{C}}B_{i}^{C}$. We set $T_{C}=T_{\theta_{C},\eta_{C}}$. Then by construction, we have an $L+3\delta$-admissible family
$\{B_{i}^{C}\}_{i\in I^{*}_{C}}=\{U_{r_{C}}\}\cup\{B_{i}^{C}:i\in J_{C}\}$
@@
$F^{\prime}(x)-F^{\prime}(y)=F(x)-F(y)-(F(x)-F(y))\llcorner B+e(x)-e(y),$
F(x)F(y)F(x)-F(y) is supported in $\bigcup_{i\in I_{C}}B_{i}^{C}\subseteq\bigcup_{i\in I^{*}_{C}}B_{i}^{C}$ and
$\operatorname{spt}((F(x)-F(y))\llcorner B)+e(x)-e(y))\subseteq B\subseteq U_{r_{C}}\subseteq\bigcup_{i\in I^{*}_{C}}B_{i}^{C}.$
@@
  1. (F1)
    $F^{\prime}|_{E}$ is localized in $\{B_{i}^{E}\}_{i\in I_{E}^{*}}$.
  2. (F2)
    $F^{\prime}(x)\llcorner(\mathbb{D}^{3}\setminus U_{r_{E}})=F(x)\llcorner(\mathbb{D}^{3}\setminus U_{r_{E}})$.
  3. (F3)
@@
  4. (F4)
    $B_{i}^{E}\cap T_{C}=\emptyset$ for every $i\in J_{E}$ and every $C\in\operatorname{Faces}_{p}(X)$ containing $E$.
  5. (F5)
    For every $x\in E$,
@@
Denote $U_{E}=U_{r_{E}}$ and $C_{E}=C_{r_{E}}$. Observe that $L+\delta\leq r_{E}\leq L+2\delta$ and that if $C_{0}\in\operatorname{Faces}_{p}(X)$ is such that $r_{E}=r_{C_{0}}$, then $C_{E}$ does not intersect $B_{i}^{C_{0}}$ for every $i\in I_{C_{0}}$ and hence neither intersects $B_{i}^{E}$ for any $i\in I_{E}$ (by monotonously $\delta$-localization). In addition, $C_{E}\cap\mathcal{S}=\emptyset$. Define $J_{E}$ and $\{B_{i}^{E}\}_{i\in I_{E}^{*}}$ as in the inductive property, so that
$\bigcup_{i\in I_{E}}B_{i}^{E}\subseteq\bigcup_{i\in I_{E}^{*}}B_{i}^{E}.$
@@
$F^{\prime}(t)=F(t)\llcorner(\mathbb{D}^{3}\setminus U_{E})+F_{E}(t)$
where we define $F_{E}$ as follows. As $\operatorname{\mathbf{M}}(F^{\prime}(x)\llcorner U_{E})$ and $\operatorname{\mathbf{M}}(F^{\prime}(y)\llcorner U_{E})$ have the same parity, we can homotop both $F^{\prime}(x)\llcorner U_{E}$ and $F^{\prime}(y)\llcorner U_{E}$ to $e(x)$ defined as 0 if $\operatorname{\mathbf{M}}(F^{\prime}(x)\llcorner U_{E})$ is even and $e$ otherwise. To be precise, for $t\in[0,\frac{1}{2}]$ we construct $F_{E}(t)$ by moving one point $q\in\operatorname{spt}(F^{\prime}(x)\llcorner U_{E})$ towards $e$ at a time. The trajectories $\gamma_{q}$ of such points are curves in $\mathbb{D}^{3}$ which do not intersect
@@
and which are supported in $\mathbb{S}^{2}=\partial\mathbb{D}^{3}$ for $q\in\mathbb{S}^{2}$. For $t=[\frac{1}{2},1]$ we proceed in the same way but starting from $F_{E}(1)=F^{\prime}(y)\llcorner U_{E}$. By construction, it holds
@@
This yields (F1) to (F5).
Inductive step. Let $E$ be a $(k+1)$-face of $X$. Choose $r_{E}\in[L+(k+1)\delta,L+(k+2)\delta]$ so that $C_{r_{E}}\cap B_{i}^{E}=\emptyset$ and $C_{r_{E}}\cap\mathcal{S}=\emptyset$ for every $i\in I_{E}$. Define $J_{E}$ and $\{B_{i}^{E}\}_{i\in I_{E}^{*}}$ as in the inductive property. Denote $U_{E}=U_{r_{E}}$, $B_{E}=B_{r_{E}}$ and $C_{E}=C_{r_{E}}$. Notice that $F^{\prime}|_{\partial E}$ is localized in $\{B_{i}^{E}\}_{i\in I_{E}^{*}}$ because of the monotonously $\delta$-localization of $F$ and the fact that